OpenStudy (anonymous):

How would i do the following : 5+7(x+1)=12+10x

OpenStudy (anonymous):

first distribute the 7, so it's 5+7x+7=12+10x then simplify the left side to 12+7x=12+10x subtract 7x from both sides: 12=12 +3x Subtract 12 from both sides: 0=3x so x=0

OpenStudy (anonymous):

5+7(x+1)=12+10x 5+7x+7=12+10x 12+7x=12+10x 12=12+10x-7x 12=12-3x 0=3x 0=x

OpenStudy (anonymous):

The answer is 0, What if i leave it as 3, would that be wrong?

OpenStudy (anonymous):

oh yes, absolutely! it's zero, that's very different from 3!

OpenStudy (anonymous):

How did you get O or 3 bye itself

OpenStudy (anonymous):

0=3x, so then divide both sides by 3. you get 0=x

OpenStudy (anonymous):

and if you want to check if 3 or 0 is the right answer, plug it into the equation, and see if you get the same number on both sides